OpenStudy (anonymous):

each of the numbers from 1 to 50 is written on a tile and the tiles are placed upside down on the top of the table. if a tile is picked up at random, what is the probability that the number on the tile is multiple of 7 or a multiple of 8?

OpenStudy (anonymous):

i get 13/50 by counting

OpenStudy (anonymous):

probably wrong though because i don't know my times tables.

OpenStudy (anonymous):

7C1/50C1+6C1/50C1

OpenStudy (anonymous):

???

OpenStudy (anonymous):

7, 14, 21, 28, 35, 42, 49 are the multiples of 7 8, 16, 24, 42, 40, 48 are the multiples of 8. i get a total of 13 so it is \[\frac{13}{50}\]

OpenStudy (amistre64):

42 is doubled

OpenStudy (amistre64):

.....well, 42 = 32 tho ;)

OpenStudy (anonymous):

well sat i also got ur answer only 7/50+6/50=13/50

OpenStudy (amistre64):

the addition rule yes; P(AorB) = P(A) + P(B) - P(AandB)

OpenStudy (anonymous):

yeah brackett i saw that you were right after i put my answer. i did think it was an odd method howewver

OpenStudy (anonymous):

yeah, for "42" read "32" you would have to look for double entries but 8*7 is bigger than 50